OpenStudy (anonymous):

Need help with integers I know the answer is 66 can someone tell me how I get to 66? I thought I do PEMDAS? 22+(-8)+32-(-12)+8

OpenStudy (anonymous):

66

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

I know the answer But I want to know how to get the answer....

OpenStudy (anonymous):

1.22+(-8)=14 2.14+36=46 3.46-(-12)=58 4.58+8=66 Answer=66

OpenStudy (anonymous):

Did you get it?

OpenStudy (tkhunny):

PEMDAS is not a thing to "do". It is an order and philosophy. Perhaps this is where you are struggling. The Order of Operations and fundamental rules of Algebra are what is important.

OpenStudy (kropot72):

The first step is to remove the brackets. It might help if the expression is rewritten in this way: 22+1(-8)+32-1(-12)+8 When the brackets are removed, you must multiply the number inside the brackets by the number outside the brackets, making sure you observe the signs of both numbers.

OpenStudy (anonymous):

Sorry but where did the 1.22 come from?I'm sorry I'm trying to refresh my memory on pre-algebra

OpenStudy (anonymous):

Aha

OpenStudy (tkhunny):

Those are line numbers. Ignore 1., 2., 3., and 4.

OpenStudy (anonymous):

ohhhhhhhhhh

OpenStudy (anonymous):

wow I feel stupid

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@kropot72 where did you get those numbers from

OpenStudy (kropot72):

@thatguy123 Are you meaning the 1 that I added after the sign before each of the brackets? 22+1(-8)+32-1(-12)+8

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

By convention there is understood to be a 1 in that position, but it is usually omitted. 22+1(-8)+32-1(-12)+8 = 22 + (1 * -8) + 32 + (-1 * -12) + 8 = 22 -8 +32 +12 + 8

OpenStudy (anonymous):

so I always should do that

OpenStudy (kropot72):

Not really. The operation that I have written should be done mentally.